The sum of three consecutive integers is 189. what are the integers?
Found 2 solutions by nyc_function, stanbon:
Answer by nyc_function(2741)

x + x + 1 + x + 2 = 169

Can you finish?

Answer by stanbon(75887)

The sum of three consecutive integers is 189. what are the integers?
----
1st: x-1
2nd: x
3rd: x+1
-----
Equation:
sum = 189
3x = 189
x = 63
----
1st: x-1 = 62
2nd: x = 63
3rd: x+1 = 64
======================
Cheers,
Stan H.